Part A With what minimum speed must you toss a 200 g ball straight up to hit the 15-m-high roof of the gymnasium if you release the ball 1.5 m above the ground? Solve this problem using energy. Express your answer using two significant figures. (answer v_min = m/s)

Part B With what speed does the ball hit the ground?

Express your answer using two significant figures. (answer v = m/s)
Answer #1

PART A)

By energy conservation we will have

mgh_1 + \frac{1}{2}mv^2 = mgH

0.200*9.8*1.5 + \frac{1}{2}*0.200*v^2 =0.200*9.8*15

v = 16.27 m/s

PART 2)

Again by energy conservation

\frac{1}{2}mv^2 + mgh_1 = \frac{1}{2} mv^2

\frac{1}{2}*0.200*16.27^2 + 0.200*9.8*1.5= \frac{1}{2}*0.200*v^2

v = 17.15 m/s